A 1,000 kg car rounds a curve on a flat road of radius 50 meters at a speed of 50km/hour. Will the car follow the curve or will it skid? Assume a) the pavement is dry and the coefficient of static friction is u=0.60; b) the pavement is icy, coefficient of friction is u=0.25?



1
Expert's answer
2022-02-20T08:19:56-0500

Given:

"R=50\:\rm m"

"v=50\:\rm km/hr=13.9\: m/s"


The maximum speed of the car to follow the curve satisfies an equation

"\frac{mv_{\max}^2}{R}=\mu mg"

or

"v_{\max}=\sqrt{\mu g R}"

(a)

"v_{\max}=\sqrt{0.60*9.8*50}=17.1\:\rm m/s>13.9\: m/s"

Hence, the car will follow the curve.

(b)

"v_{\max}=\sqrt{0.25*9.8*50}=11.1\:\rm m/s<13.9\: m/s"

Hence, the car will skid the curve.